Christopher's Cranks uses a machine that can produce 100 cranks per hour. The firm operates 12 hours per day, five days per week. Due to regularly scheduled preventive maintenance, the firm expects the machine to be running during approximately 95% of the available time. Based on experience with other products, the firm expects to achieve an efficiency level for the cranks of 85%. What is the expected weekly output of cranks for this company?
Answer #1

Answer:

Ideally:

  • Production per hour = 100 crank
  • Production per day = hours per day * production per hour =12*100 = 1200 crank
  • Production per week =Days per week * production per day = 5 * 1200 = 6000 cranks

Maintenance adjusted:

  • Since output will 95% of what it would be ideally due to maintenance.
  • Maintenance adjusted output = 0.95*6000 = 5700 units

Efficiency adjusted:

  • Also efficiency level is to be 85 %, hence actual output will be 85 % of maintenance adjusted output.
  • Actual output = 0.85*5700 = 4,845 cranks (answer)